OpenStudy (anonymous):

A building has an entry the shape of a parabolic arch 84 ft high and 42 ft wide at the base as shown below. Find an equation for the parabola if the vertex is put at the origin of the coordinate system.

hero (hero):

"as shown below" implies that a picture is given. Please post it.

OpenStudy (anonymous):

are you like supposed to solve it like this 21^2=4p(-84) 441/-84=4p -5.25=4p or x^2=-5.3y

OpenStudy (anonymous):

okay...I got that part

hero (hero):

We know it is a parabola so try using this formula y = a(x - h)^2 + k You already know the vertex (0,0) and two points on the parabola (21, -84), (-21,-84)

hero (hero):

Sorry, sign errors

hero (hero):

You need to make two equations, one for each given point, then solve for a

hero (hero):

-84 = a(21 - 0)^2 + 0 -84 = a(-21 - 0)^2 + 0 -84 = a(21)^2 -84 = a(-21)^2 -84 = 441a -84/441 = a

hero (hero):

-4/21 = a

OpenStudy (anonymous):

my answers are supposed to be like this y2 = -21x x2 = -21y x2 = -5.3y y2 = -5.3x

hero (hero):

Hmmm

OpenStudy (anonymous):

I thought that because the parabola opened down it was supposed to look like x^2= somthing

hero (hero):

It is \[y = -\frac{4}{21}x^2\]

hero (hero):

That's the parabola I got

hero (hero):

If you graph that, you will see that (-21, -84) and (21, 84) are points on it

hero (hero):

I think I see. We have to put the parabola in a different form that matches one of the answer choices.

hero (hero):

Try isolating x^2

OpenStudy (anonymous):

yeah that's what I just did

OpenStudy (anonymous):

but, why did you use that formula ?

OpenStudy (anonymous):

never mind ...would you mind helping me with another problem ?Find an equation in standard form for the hyperbola with vertices at (0, ±8) and asymptotes at .

hero (hero):

Because that formula is the equation for a parabola when you know at least one point on the parabola and the vertex. It allows you to plug in two points into one equation.

OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±8) and asymptotes at .

hero (hero):

Use the formula for equation of a hyperbola to help you

OpenStudy (anonymous):

okay I tried doing this a=1*2=2^2=4 b=2*2=4^2=16 y^2/64-x^2/16

OpenStudy (anonymous):

is that right?

hero (hero):

Oh, sorry, I didn't see that you posted a response until now.